OpenStudy (anonymous):

How does the division rule for exponents help in understanding why anything to the zero power is 1?

OpenStudy (anonymous):

do u still need some help here

OpenStudy (anonymous):

WELL \[x \neq0\] \[x^1=x\] \[x^{-1}=1/x\] \[x^{-1}*x^1 = x^{-1+1}= x^0=x*\frac{1}{x}=\frac{x}{x}=\frac{1}{1}=1\]

OpenStudy (anonymous):

ok iam a little cufused on this question

OpenStudy (anonymous):

do u have mre info to give me here to bettre help ou here

OpenStudy (anonymous):

\[\frac{1}{x^n}=x^{-n}\]

OpenStudy (anonymous):

Boom!! Mind blown. Good work @completeidiot.